# Non-Collaborative User Simulators for Tool Agents

### v3

@@ B.3 Ending Verifier
User: I’$l$ add one checked bag, and I don’t need travel insurance this time. Please use my credit card for payment.
Agent: Please confirm the following booking details: Flight HAT063 on 2024-05-29 from Dallas to Newark in economy class for Olivia Gonzalez with one checked bag and no travel insurance.
Payment will be made using your credit card ending in 9475. Confirm with ’yes’ to proceed or let me know if you want any changes.
@@ C.1 Unavailable Service
We provide the examples of unavailable user goal generated during our experiments. These additional goals are attached to the original user goal with the sentence “In addition to the above conditions, if the agent is able to fulfill them, also include the conditions from {unavailable_user_goal_list}.” as in Figure 3-a.
Table 5: Examples of unavailable user goal
MultiWOZ
“You want to know the menu details and nutritional information for the thai restaurant before booking a table.”,
“You want to request a room with a specific view, such as a garden or city view, in the hotel booking.”,
“You want a train that has wheelchair accessibility features.”
$\tau$-bench
“You want to arrange a special meal preference for the cheapest economy flight you plan to book.”
@@ C.2 Tangential
We define four tangential dialogue acts motivated by open-domain dialogue studies midas_tangential as follows:
@@ C.3 Impatience
We define three impatient dialogue acts referring bellgerent_abuse and threat as follows:
@@ D.1 Evaluation Environment Implementation Details
We used the database from the MultiWOZ 2.4 dataset repository as is, and implemented the API ourselves in Python. We designed each domain to have one retrieve API and one book API, except for the taxi domain which only has a book API. We also included a “book_cancel” API. Additionally, we provided helper APIs from appworld, allowing the tool agent to utilize a total of 11 APIs during dialogue simulation. The tool agent uses the helper API as shown in Figure 6. First, it queries the available app (domain) descriptions using the “show_app_description” API. Then, it selects the appropriate app based on the context and queries the API description using “show_api_description”. Finally, the full documentation of the selected API is retrieved through “show_api_documentation”. In MultiWOZ, we commonly give all tool agents the following instruction prompt shown in Listing 17.
@@
We use the existing work’s environment with almost no modifications, and the only difference is the removal of the “transfer_to_human” API. Full implementation such as APIs, database and tool agent prompts of $\tau$-bench are in the Appendix of tau_bench
@@ D.3 Goal Alignment Metric
To measure the reliability of the judge model for goal alignment, we sampled 40 dialogues from Tau-bench and MultiWOZ, including both collaborative and non-collaborative user simulations (20 goal-aligned dialogues and 20 goal-misaligned dialogues), and instructed a human annotator (graduate student) to evaluate whether all information specified in the user goal was provided, given the user goal and dialogue. We then measured the Matthews correlation coefficient (MCC) between the human judgments and the judge model’s judgments (the MCC is the Pearson correlation coefficient for binary annotations). The correlation was 0.77, which is generally considered strong (mcc_strong) . This result suggests that the judge model has strong agreement with human judgments.
@@ D.7 Fine-tuning Non-collaborative User Data
Performance
Table 20 shows that the uniformly weighted approach achieved overall performance gains across all behaviors compared to training only on collaborative dialogues (see the “Only Collaborative”), with an average performance of 86.9%. However, this setup led to a relatively low score on Incomplete Utterance. For the non-uniformly weighted approach, which includes more training data for Incomplete Utterance, we observed performance improvement in the Incomplete Utterance mode. The average performance remains nearly identical at 86.6%.
Only Collaborative 91.6 61.2 83.1 85.1 73.0 78.8
Uniformly weighted 93.5 85.7 87.4 89.6 78.4 86.9
Non-uniformly weighted 91.6 85.7 85.7 87.6 82.3 86.6
Practical Deployment
Based on our experimental results and the characteristics of real-world industry services (opinion_leader), uniform weighting of customer types seems to be a reasonable choice for maintaining overall service satisfaction and agent performance. However, if there are specific user types that require more attention in certain domains, or if there exist minimum performance criteria across all types of customers, it is reasonable to use more training data for those categories.
@@ D.8 Non-Collaborative User Behavior co-occurrence Patterns
Unavailable service requests occur when users are unaware of the agent’s capabilities, impatience arises when users become frustrated due to services not working as expected, tangential behavior emerges when users seek sustained attention, and incomplete utterances can appear due to humans’ principle of least effort (least_effort) or mistakes in typing situations. Since these behaviors originate from different root causes and motives, they can plausibly appear in combination, though certain combinations may be less natural. The explanation is as follows.
@@
Requests for unavailable service can naturally co-occur with all the other behaviors. Specifically, Unavailable Service + Tangential can occur when tangential users are unaware of agent capabilities, and Unavailable Service + Incomplete Utterance can arise in typing environments when users don’t know the chatbot’s capabilities. Unavailable Service + Impatience has been characterized in marketing research as customers with a sense of entitlement who believe that employees should do anything for them (non_coll_unavailable; entitlement). Similarly, Incomplete Utterance + Impatience can be defined as users who interact with online chatbots through typing and become frustrated when the chatbot fails to address their requests (crolic2022blame). The other two combinations seem less natural. For Incomplete Utterance + Tangential, incomplete utterances arise in typed communication, whereas tangential utterances are more likely in online posts or face-to-face interactions. Hence, the combination of these behaviors may be relatively infrequent. In the case of Impatience + Tangential, impatience is a behavior arising from users wanting fast service while tangential behavior delays service progress, making the simultaneous performance of both behaviors less likely.
@@ E.1 Combined Behavior Evaluation
We randomly sampled 30 test scenarios from the combined pool of MultiWOZ (89 scenarios) and $\tau$-bench (157 scenarios) using a fixed random seed of 42 to ensure reproducibility. This sample size was chosen to balance evaluation comprehensiveness with annotator workload, while maintaining statistical reliability across six different behavior combinations.
Evaluation Protocol
Three independent annotators evaluated each dialogue pair through pairwise comparison. For each of the 30 sampled scenarios, we presented dialogue pairs generated by our simulator and PBUS baseline, both exhibiting the same paired non-collaborative behaviors. To eliminate position bias, the presentation order of dialogues (A or B) was randomized for each pair. The behavior combinations evaluated were: (1) Impatience + Unavailable Service, (2) Tangential + Incomplete Utterance, (3) Tangential + Unavailable Service, (4) Impatience + Tangential, (5) Incomplete Utterance + Unavailable Service, and (6) Incomplete Utterance + Impatience.
Evaluation Dimensions
Annotators assessed three dimensions for each dialogue pair:
  • First Behavior Realism: Which dialogue more realistically simulates the first non-collaborative behavior (e.g., “Which user’s attempt to go off-topic feels more realistic?”)
  • Second Behavior Realism: Which dialogue more realistically simulates the second non-collaborative behavior (e.g., “Which user’s expression of impatience is more realistic and human-like?”)
  • Overall Coherence: Which dialogue presents a more natural and coherent combination of both behaviors
E.2 Single Behavior Evaluation
Sampling Methodology
Following the same sampling procedure as the combined behavior evaluation, we randomly sampled 30 test scenarios.
@@ E.2 Single Behavior Evaluation
Unlike the combined behavior evaluation, annotators assessed a single dimension:
@@ F.2 Performance
Collaborative Unavailable Tangential Impatience Incomplete Utterance
@@
Collaborative Unavailable Tangential Impatience Incomplete Utterance
@@
Collaborative Unavailable Tangential Impatience Incomplete Utterance
@@
Collaborative Unavailable Tangential Impatience Incomplete Utterance
@@
Analysis
Similar to Table 1, GPT-4.1-mini demonstrates robustness to non-collaborative behaviors, while other models show overall performance degradation under non-collaborative conditions. For ColBench, which involves programming tasks, we observed performance drops in the tangential and incomplete utterance modes, similar to the trends in MultiWOZ and $\tau$-bench. In contrast, MINT, where the user provides feedback on collaboration tasks, did not exhibit such trends. Additionally, while Qwen3-30b-a3b showed robustness to unavailable service requests in ColBench, MultiWOZ, and $\tau$-bench, it demonstrated performance degradation in MINT. This suggests that domains aimed at fulfilling the user simulator’s goal exhibit similar model performance patterns even in non-booking contexts. However, domains with fundamentally different characteristics, such as solving goals through collaboration with the user simulator, may show different model performance patterns.
@@ G.2 Tool Agent Prompts
We provide the MultiWOZ tool agent prompt we created. The tool agent prompt for $\tau$-bench can be found in the appendix of tau_bench.

### v5

@@ B.3 Ending Verifier
User: I’ll add one checked bag, and I don’t need travel insurance this time. Please use my credit card for payment.
Agent: Please confirm the following booking details: Flight HAT063 on 2024-05-29 from Dallas to Newark in economy class for Olivia Gonzalez with one checked bag and no travel insurance. Payment will be made using your credit card ending in 9475. Confirm with ‘yes’ to proceed or let me know if you want any changes.
@@ C.1 Unavailable Service
We provide the examples of unavailable user goal generated during our experiments in Table  7. These additional goals are attached to the original user goal with the sentence “In addition to the above conditions, if the agent is able to fulfill them, also include the conditions from {unavailable_user_goal_list}.” as in Figure 3-a.
@@ C.2 Tangential
We define four tangential dialogue acts motivated by open-domain dialogue studies Yu and Yu (2021) as follows:
@@
Table 7: Examples of unavailable user goal
MultiWOZ “You want to know the menu details and nutritional information for the thai restaurant before booking a table.”
“You want to request a room with a specific view, such as a garden or city view, in the hotel booking.”
“You want a train that has wheelchair accessibility features.”
$\tau$-bench “You want to arrange a special meal preference for the cheapest economy flight you plan to book.”
@@ C.3 Impatience
We define three impatient dialogue acts referring Reynolds and Harris (2009) and Harris and Reynolds (2004) as follows:
@@ D.1 Evaluation Environment Implementation Details
We used the database from the MultiWOZ 2.4 dataset repository as is, and implemented the API ourselves in Python. We designed each domain to have one retrieve API and one book API, except for the taxi domain which only has a book API. We also included a “book_cancel” API. Additionally, we provided helper APIs from Trivedi et al. (2024), allowing the tool agent to utilize a total of 11 APIs during dialogue simulation (see the Table  9). The tool agent uses the helper API as shown in Figure 6. First, it queries the available app (domain) descriptions using the “show_app_description” API. Then, it selects the appropriate app based on the context and queries the API description using “show_api_description”. Finally, the full documentation of the selected API is retrieved through “show_api_documentation”. In MultiWOZ, we commonly give all tool agents the following instruction prompt shown in Listing 17.
@@
We use the existing work’s environment with almost no modifications, and the only difference is the removal of the “transfer_to_human” API. Full implementation such as APIs, database and tool agent prompts of $\tau$-bench are in the Appendix of Yao et al. (2024)
@@ D.3 Goal Alignment Metric
To measure the reliability of the judge model for goal alignment, we sampled 40 dialogues from Tau-bench and MultiWOZ, including both collaborative and non-collaborative user simulations (20 goal-aligned dialogues and 20 goal-misaligned dialogues), and instructed a human annotator (graduate student) to evaluate whether all information specified in the user goal was provided, given the user goal and dialogue. We then measured the Matthews correlation coefficient (MCC) between the human judgments and the judge model’s judgments (the MCC is the Pearson correlation coefficient for binary annotations). The correlation was 0.77, which is generally considered strong (Chicco and Jurman, 2020) . This result suggests that the judge model has strong agreement with human judgments.
@@ D.8 Non-Collaborative User Behavior co-occurrence Patterns
Unavailable service requests occur when users are unaware of the agent’s capabilities, impatience arises when users become frustrated due to services not working as expected, tangential behavior emerges when users seek sustained attention, and incomplete utterances can appear due to humans’ principle of least effort (Zipf, 2016) or mistakes in typing situations. Since these behaviors originate from different root causes and motives, they can plausibly appear in combination, though certain combinations may be less natural. The explanation is as follows.
@@
Requests for unavailable service can naturally co-occur with all the other behaviors. Specifically, Unavailable Service + Tangential can occur when tangential users are unaware of agent capabilities, and Unavailable Service + Incomplete Utterance can arise in typing environments when users don’t know the chatbot’s capabilities. Unavailable Service + Impatience has been characterized in marketing research as customers with a sense of entitlement who believe that employees should do anything for them (Bitner et al., 1990; Boyd III and Helms, 2005). Similarly, Incomplete Utterance + Impatience can be defined as users who interact with online chatbots through typing and become frustrated when the chatbot fails to address their requests (Crolic et al., 2022). The other two combinations seem less natural. For Incomplete Utterance + Tangential, incomplete utterances arise in typed communication, whereas tangential utterances are more likely in online posts or face-to-face interactions. Hence, the combination of these behaviors may be relatively infrequent. In the case of Impatience + Tangential, impatience is a behavior arising from users wanting fast service while tangential behavior delays service progress, making the simultaneous performance of both behaviors less likely.
@@ E.1 Evaluation
We randomly sampled 30 test scenarios from the combined pool of MultiWOZ (89 scenarios) and $\tau$-bench (157 scenarios). This sample size was chosen to balance evaluation comprehensiveness with annotator workload, while maintaining statistical reliability across six different behavior combinations.
@@
Annotators assessed a following question for each dialogue pair:
@@ G.2 Tool Agent Prompts
We provide the MultiWOZ tool agent prompt we created. The tool agent prompt for $\tau$-bench can be found in the appendix of Yao et al. (2024).